# Error-Minimizing Measurements in Postselected One-Shot Symmetric Quantum State Discrimination and Acceptance as a Performance Metric

### v1

@@ III-A The set of all error minimizing measurements and an arbitrary construction
$\rho=\frac{3}{4}|+\rangle\langle+|+\frac{1}{4}|-\rangle\langle-|=\frac{1}{2}|0\rangle\langle 0|+\frac{1}{4}|1\rangle\langle 0|+\frac{1}{4}|0\rangle\langle 1|+\frac{1}{2}|1\rangle\langle 1|,\qquad\sigma=\frac{3}{4}|0\rangle\langle 0|+\frac{1}{4}|1\rangle\langle 1|.$
Note that, in this example $\Pi_{\rho}=\Pi_{\sigma}=\mathrm{I}$. We obtain
@@
We get eigenvalues of $\sigma^{-1/2}\rho\sigma^{-1/2}$ as $\frac{4\pm\sqrt{7}}{3}$ and corresponding eigenvectors as $\frac{{\sqrt{3}|0\rangle+(2\pm\sqrt{7})|1\rangle}}{\sqrt{14\pm 4\sqrt{7}}}.$ Thus $R_{\max}(\rho,\sigma)=\frac{4+\sqrt{7}}{3},R_{\min}(\rho,\sigma)=\frac{4-\sqrt{7}}{3},R_{\max}(\sigma,\rho)=\frac{4+\sqrt{7}}{3}.$ So, we get
@@
As we obtained $p_{\rho}^{*}=1/2$ and $p_{\sigma}^{*}=1/2$, using Theorem III.1, we obtain minimum error and error minimizing condition as the following.
@@
Note that $\mathrm{T}^{\max}$ and $\mathrm{T}^{\min}$ are both rank 1 and so, any element in the set $\mathcal{P}(\mathrm{T}^{\max})$ and $\mathcal{P}(\mathrm{T}^{\max})$ are of the form $c\mathrm{T}^{\max}$ and $c\mathrm{T}^{\min}$ for some $c\geq 0$. Following the fact that $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$ have simple form, and that $\Pi_{\sigma}=\mathrm{I}$ in this example, we can obtain constraint on $\Lambda_{\rho},\Lambda_{\sigma}$ by simply multiplying $\sigma^{-1/2}$ on both sides, a more general way to find most general form of $\Lambda_{\rho},\Lambda_{\sigma}$ for error-minimization is given next.
Theorem III.1 gives the subspace where $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$ lie for $\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}$ to be an error minimizing measurement. Building on Theorem III.1, we now focus on deriving the set of measurements that achieve minimum postselected symmetric error. We use Lemma C.1 from Appendix C to obtain the constraint on $\Lambda_{\rho}$ and $\Lambda_{\sigma}$, from constraint on $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$ as given in Theorem III.1 to ensure that the measurement achieves the minimum postselected asymmetric error. Then, we parameterize $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ in a way to satisfy the obtained constraints and further put conditions $\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$ to ensure that it remains a valid measurement, thus obtaining the set of all measurements. The error minimizing set and relevant proof is stated in the following theorem.
Theorem III.2.
@@ 3. $\underline{\mathcal{C}\mathrm{3}}$:
Here $\mathrm{P}^{\max}=\Pi_{\sigma^{-1/2}\mathrm{T}^{\max}\sigma^{-1/2}}$ and $\mathrm{P}^{\min}=\Pi_{\sigma^{-1/2}\mathrm{T}^{\min}\sigma^{-1/2}}$ with $\mathrm{T}^{\max}\stackrel{{\scriptstyle\Delta}}{{=}}\Pi^{\max}_{\sigma^{-1/2}\rho\sigma^{-1/2}}$ and $\mathrm{T}^{\min}\stackrel{{\scriptstyle\Delta}}{{=}}\Pi^{\min}_{\sigma^{-1/2}\rho\sigma^{-1/2}}$ as defined in Theorem III.1.
@@
Using Theorem C.1 and Lemma C.1 from Appendix C, we get
@@
We will use these two results for $\Gamma=\Lambda_{\rho}$ and $\Gamma=\Lambda_{\sigma}$ to get the expression for $\Lambda_{\rho}$ and $\Lambda_{\sigma}$. We begin with condition on $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ derived in Theorem III.1 for the three cases separately.
@@ 1. $\underline{\mathcal{C}\mathrm{1}}$
    In this case, using Theorem III.1, for error minimizing POVM, we get the condition $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}=0$. Now, using equivalence shown in (22), we get
@@
    A general $\Lambda_{\rho}$ can be chosen as $\Lambda_{\rho}={c}\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}$ for some $\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max}),c\geq 0$. On putting the condition that $\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$, we obtain $\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho}$, and for any such $\Lambda_{\sigma}$ to exist, it has to be ensured that $\Lambda_{\rho}\leq\mathrm{I}$ and thus $c\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}\leq\mathrm{I}\Rightarrow c\leq\left\|\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}\right\|_{\infty}^{-1}$. Writing these condition together, we get an error minimizing measurement as
@@ 2. $\underline{\mathcal{C}\mathrm{2}}$
    In this case, using Theorem III.1, for error minimizing POVM, we get the condition $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}=0,\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min})$. Now, using equivalence shown in (22), we get
@@
    A general $\Lambda_{\sigma}$ can be chosen as $\Lambda_{\sigma}={c}\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}$ for some $\psi_{\min}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\min}),c\geq 0$. On putting the condition that $\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$, we obtain $\Lambda_{\rho}\leq\mathrm{I}-\Lambda_{\sigma}$, and for any such $\Lambda_{\rho}$ to exist, it has to be ensured that $\Lambda_{\sigma}\leq\mathrm{I}$ and thus $c\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\leq\mathrm{I}\Rightarrow c\leq\left\|\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\right\|_{\infty}^{-1}$. Writing these condition together, we get an error minimizing measurement as
@@ 3. $\underline{\mathcal{C}\mathrm{3}}$
    In this case, using Theorem III.1, for error minimizing POVM, we get the condition $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min})$. Now, using equivalence shown in (22), we get
@@
    Writing this as set, we obtain the result stated in the theorem.
Constructing an arbitrary error minimizing measurement: Theorem III.2 gives an intuitive way to construct any POVM that minimizes error. The way to choose parameters for three cases are considered separately as given in the table below. The first column shows the parameters and rest three show the way to choose them in different cases.
@@
TABLE I: Table showing a parameterization of an arbitrary error minimizing measurement
The measurement POVM is given as $\{\Lambda_{\sigma},\Lambda_{\rho},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}$. This is the most general way to construct an arbitrary POVM achieving the minimum postselected symmetric error. This way of constructing the set also gives us more freedom in choosing the measurement in the sense that any possible set of measurements can be obtained by choosing appropriate set for the free variables $\psi_{\max},\psi_{\min},c$ and $c_{r}$.
Remark 4.
For unique characterization of an arbitrary POVM in the set $\mathcal{E}_{s}(\rho,\sigma,p)$, we have:
  1. For the case $\underline{\mathcal{C}\mathrm{1}}$, the set of parameters
    $(c,\psi_{\max},\Lambda_{\sigma})$ is sufficient.
  2. For the case $\underline{\mathcal{C}\mathrm{2}}$, the set of parameters
    $(c,\psi_{\min},\Lambda_{\rho})$ is sufficient.
  3. For the case $\underline{\mathcal{C}\mathrm{3}}$, the set of parameters
    $(c,c_{r},\psi_{\max},\psi_{\min})$ is sufficient.
The above set of values give a parameterization of an arbitrary error minimizing measurement. We will utilize this later when we calculate acceptance in terms of these parameters in a form which can be easily maximized. Interestingly, as we will see, the acceptance turns out to be only a function of $c$ and $c_{r}$.
Finding set of error minimizing measurements for system given in the Example 2(a) is illustrated below. In this example, we will see that maximum value of $c$ can be obtained easily, which can be later utilized to find the maximum acceptance.
@@ 3. For the case $\underline{\mathcal{C}\mathrm{3}}$, the set of parameters
Referring to system described in Example 2(a), we get
@@
For error minimizing measurement, we have to take $\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max})$ and $\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max})$. Note that, $\Pi_{\sigma}=\mathrm{I}$ and $\mathrm{P}^{\max}$ and $\mathrm{P}^{\min}$ are rank 1, and so we get $\psi_{\max}=\mathrm{P}^{\max}$ and $\psi_{\min}=\mathrm{P}^{\min}$. Further $\mathrm{Tr}\left(\psi_{\max}\sigma\right)=\mathrm{Tr}\left(\mathrm{P}^{\max}\sigma\right)=\frac{1}{4}\frac{14+4\sqrt{7}}{12+4\sqrt{7}}$ and $\mathrm{Tr}\left(\psi_{\min}\sigma\right)=\mathrm{Tr}\left(\mathrm{P}_{\min}\sigma\right)=\frac{1}{4}\frac{14-4\sqrt{7}}{12-4\sqrt{7}}$. So, most general error minimizing measurement is given as below.
@@
    If $p_{\rho}>1/2:$ maximum value of c is $\frac{1}{4}\frac{14+4\sqrt{7}}{12+4\sqrt{7}}=0.272$.
@@
    If $p_{\rho}<1/2:$ maximum value of c is $\frac{1}{4}\frac{14-4\sqrt{7}}{12-4\sqrt{7}}=0.603$.
@@
    If $p_{\rho}=1/2:$ maximum value of c is $\max_{0\leq c_{r}\leq 1}\left\|c_{r}\frac{4(12+4\sqrt{7})}{14+4\sqrt{7}}\mathrm{P}^{\max}-(1-c_{r})\frac{4(12-4\sqrt{7})}{14-4\sqrt{7}}\mathrm{P}^{\min}\right\|_{\infty}^{-1}$.
Here, we are obtaining a simple numerical bound on $c$ in first two cases because $\mathrm{P}^{\max}$ and $\mathrm{P}^{\min}$ are rank 1 operators. Usually, it depends on $\psi_{\max}$ and $\psi_{\min}$. We will see in the next subsection that acceptance depends only on $c$ except for the third case. So, easily obtained maximum value of $c$ helps in finding maximum acceptance.
III-B The maximum acceptance for error-minimizing measurements
Note that Theorem III.2 gives the set of measurements that minimize the postselected symmetric error as $\mathcal{E}_{s}(\rho,\sigma,p)$ and gives their parameterization in terms of $(c,\psi_{\max},\Lambda_{\sigma})$, $(c,\psi_{\min},\Lambda_{\rho})$ or $(c,c_{r},\psi_{\max},\psi_{\min})$. The next Lemma gives acceptance for a given error minimizing measurement as a function of parameters $c$ and $c_{r}$.
Theorem III.3.
For the most general $\Lambda\in\mathcal{E}_{s}(\rho,\sigma,p)$, as given in Theorem III.2, the acceptances when the unknown states $\rho$ or $\sigma$, are respectively given below.

### v2

@@ 3. $\bullet$ Case $\underline{\mathcal{C}\mathrm{3}}$:
$\rho=\frac{3}{4}|+\rangle\langle+|+\frac{1}{4}|-\rangle\langle-|=\frac{1}{2}|0\rangle\langle 0|+\frac{1}{4}|1\rangle\langle 0|+\frac{1}{4}|0\rangle\langle 1|+\frac{1}{2}|1\rangle\langle 1|,\quad\text{and }\sigma=\frac{3}{4}|0\rangle\langle 0|+\frac{1}{4}|1\rangle\langle 1|.$
Note that, in this example, $\Pi_{\rho}=\Pi_{\sigma}=\mathrm{I}$. We obtain
@@
We get eigenvalues of $\sigma^{-1/2}\rho\sigma^{-1/2}$ as $\frac{4\pm\sqrt{7}}{3}$ and corresponding eigenvectors as $\frac{{\sqrt{3}|0\rangle+(2\pm\sqrt{7})|1\rangle}}{\sqrt{14\pm 4\sqrt{7}}}.$ Thus $R_{\max}(\rho,\sigma)=\frac{4+\sqrt{7}}{3},R_{\min}(\rho,\sigma)=\frac{4-\sqrt{7}}{3}$. Similarly, $R_{\max}(\sigma,\rho)=\frac{4+\sqrt{7}}{3}.$ So, we get
@@
As $p_{\rho}^{*}=p_{\sigma}^{*}=1/2$, using Theorem 1, we obtain the minimum error and error minimizing condition as the following.
@@
Note that $\mathrm{T}^{\max}$ and $\mathrm{T}^{\min}$ are both rank 1 in this example and so, any element in the set $\mathcal{P}(\mathrm{T}^{\max})$ and $\mathcal{P}(\mathrm{T}^{\max})$ are of the form $c\mathrm{T}^{\max}$ and $c\mathrm{T}^{\min}$ for some $c\geq 0$. Following the fact that $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$ have simple form, and that $\Pi_{\sigma}=\mathrm{I}$ (thus $\sigma^{-1/2}(\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2})\sigma^{-1/2}=\Pi_{\sigma}\Lambda_{\sigma}\Pi_{\sigma}=\Lambda_{\sigma}$, and same for $\Lambda_{\rho}$) in this example, we can obtain constraint on $\Lambda_{\rho},\Lambda_{\sigma}$ by simply multiplying $\sigma^{-1/2}$ on both sides. A comprehensive way to find the most general form of $\Lambda_{\rho},\Lambda_{\sigma}$ for error-minimization is given next.
Theorem 1 gives the subspace where $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$ should lie for $\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}$ to be an error minimizing measurement. Building on Theorem 1, we now focus on deriving the set of all error-minimizing measurements. We start with the constraint on $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$ given in (11) and use Lemma B.1 from Appendix B to obtain the constraint on $\Lambda_{\rho}$ and $\Lambda_{\sigma}$, from constraint on $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$. Then, we parameterize $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ to satisfy the obtained constraints and then put condition $\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$ to ensure that it remains a valid measurement, thus obtaining the set of all measurements. The set of all error-minimizing measurements is obtained in the following theorem.
Theorem 2.
@@ 3. $\underline{\mathcal{C}\mathrm{3}}$:
Here $\mathrm{P}^{\max}=\Pi_{\sigma^{-1/2}\mathrm{T}^{\max}\sigma^{-1/2}}$ and $\mathrm{P}^{\min}=\Pi_{\sigma^{-1/2}\mathrm{T}^{\min}\sigma^{-1/2}}$ with $\mathrm{T}^{\max}\stackrel{{\scriptstyle\Delta}}{{=}}\Pi^{\max}_{\sigma^{-1/2}\rho\sigma^{-1/2}}$ and $\mathrm{T}^{\min}\stackrel{{\scriptstyle\Delta}}{{=}}\Pi^{\min}_{\sigma^{-1/2}\rho\sigma^{-1/2}}$ as defined in Theorem 1.
@@
Using Theorem B.1 and Lemma B.1 from Appendix B, we get
@@
We will use these two results for $\Gamma=\Lambda_{\rho}$ and $\Gamma=\Lambda_{\sigma}$ to get the expression for $\Lambda_{\rho}$ and $\Lambda_{\sigma}$. We begin with the condition on $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ derived in Theorem 1 for the three cases separately.
@@ 1. $\underline{\mathcal{C}\mathrm{1}}$
    In this case, using Theorem 1, for error minimizing POVM, we get the condition $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}=0$. Now, using equivalence shown in (22), we get
@@
    A general $\Lambda_{\rho}$ can be chosen as $\Lambda_{\rho}={c}\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}$ for some $\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max}),c\geq 0$. On putting the condition that $\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$, we obtain $\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho}$, and for any such $\Lambda_{\sigma}$ to exist, it has to be ensured that $\Lambda_{\rho}\leq\mathrm{I}$ and thus $c\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}\leq\mathrm{I}\Rightarrow c\leq\left\|\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}\right\|_{\infty}^{-1}$. Writing these conditions together, we get an error-minimizing measurement as
@@ 2. $\underline{\mathcal{C}\mathrm{2}}$
    In this case, using Theorem 1, for error minimizing POVM, we get the condition $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}=0,\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min})$. Now, using equivalence shown in (22), we get
@@
    A general $\Lambda_{\sigma}$ can be chosen as $\Lambda_{\sigma}={c}\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}$ for some $\psi_{\min}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\min}),c\geq 0$. On putting the condition that $\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$, we obtain $\Lambda_{\rho}\leq\mathrm{I}-\Lambda_{\sigma}$, and for any such $\Lambda_{\rho}$ to exist, it has to be ensured that $\Lambda_{\sigma}\leq\mathrm{I}$ and thus $c\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\leq\mathrm{I}\Rightarrow c\leq\left\|\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\right\|_{\infty}^{-1}$. Writing these conditions together, we get an error-minimizing measurement as
@@ 3. $\underline{\mathcal{C}\mathrm{3}}$
    In this case, using Theorem 1, for error minimizing POVM, we get the condition $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min})$. Now, using equivalence shown in (22), we get
@@
Writing this as a set, we obtain the result stated in the theorem. ∎
Remark 4.
While the work in [33] gives an example that achieves the minimum error, we have given the most general structure of error-minimizing measurements. The example of error-minimizing measurement given in [33] satisfies the condition for error-minimizing measurements given in Theorem 1 and belongs to the set of error-minimizing measurements given in Theorem 2. So, while the error obtained by example in [33] and our obtained set of measurements $\mathcal{E}_{s}$ is the same, [33] gives a single example, whereas we provide the complete set of all error-minimizing measurements.
Constructing an arbitrary error minimizing measurement: Theorem 2 gives an intuitive way to construct any error minimizing measurement, which is described in Table III. The first column shows the parameters and the following three show how to choose them for different case $\mathcal{C}1,\mathcal{C}2$, and $\mathcal{C}3$. To construct an arbitrary error minimizing measurement, we first select $\psi_{\max}$ and $\psi_{\min}$ according to the applicable case, then choose $c_{r}$ and $c$ as given in the second row and then construct $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ as given in the last row. The measurement $\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}$ thus constructed will be an error minimizing measurement.
@@
TABLE III: Parameterization of error minimizing measurements and method to construct an arbitrary error minimizing measurement.
Note that this method giving the general and parametric form of measurements, provides us with freedom in choosing the measurement by varying $\psi_{\max},\psi_{\min},c$ and $c_{r}$.
We now give another example that illustrates how to find the set of error-minimizing measurements for the system given in the Example 2(a).
@@
Applying definition of $\mathrm{P}^{\max}$ and $\mathrm{P}^{\min}$ given in Theorem 2 to the system described in Example 2(a), we get
@@
Further, from theorem 2, we have $\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max})$ and $\psi_{\min}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\min})$. Note that $\Pi_{\sigma}=\mathrm{I}$ and $\mathrm{P}^{\max}$ and $\mathrm{P}^{\min}$ are rank 1, so the only element in $\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max})=\mathcal{S}(\mathrm{P}^{\max})$ is $\mathrm{P}^{\max}$ and the only element in $\mathcal{S}(\mathrm{P}^{\min})$ is $\mathrm{P}^{\min}$. Hence, we get $\psi_{\max}=\mathrm{P}^{\max}$ and $\psi_{\min}=\mathrm{P}^{\min}$. Further $\mathrm{Tr}\left(\psi_{\max}\sigma\right)=\mathrm{Tr}\left(\mathrm{P}^{\max}\sigma\right)=\frac{1}{4}\frac{14+4\sqrt{7}}{12+4\sqrt{7}}$ and $\mathrm{Tr}\left(\psi_{\min}\sigma\right)=\mathrm{Tr}\left(\mathrm{P}_{\min}\sigma\right)=\frac{1}{4}\frac{14-4\sqrt{7}}{12-4\sqrt{7}}$. So, the set of error minimizing measurements can be given as below.
@@
    If $p_{\rho}>1/2:$ the maximum value of c is $\frac{1}{4}\frac{14+4\sqrt{7}}{12+4\sqrt{7}}=0.272$.
@@
    If $p_{\rho}<1/2:$ the maximum value of c is $\frac{1}{4}\frac{14-4\sqrt{7}}{12-4\sqrt{7}}=0.603$.
@@
    If $p_{\rho}=1/2:$ the maximum value of c is $\max_{0\leq c_{r}\leq 1}\left\|c_{r}\frac{4(12+4\sqrt{7})}{14+4\sqrt{7}}\mathrm{P}^{\max}-(1-c_{r})\frac{4(12-4\sqrt{7})}{14-4\sqrt{7}}\mathrm{P}^{\min}\right\|_{\infty}^{-1}$.
We are able to obtain a numerical value of bound on $c$ in the first two cases because $\mathrm{P}^{\max}$ and $\mathrm{P}^{\min}$ are rank 1 operators. Usually, the depends on $\psi_{\max}$ and $\psi_{\min}$. We will see in the next subsection that acceptance depends only on $c$ except for the third case. So, easily obtaining the maximum value of $c$ helps find the maximum value of acceptance.
III-B The maximum acceptance achieved by error-minimizing measurements
Note that Theorem 2 gives the set of all measurements that minimize the postselected symmetric error as $\mathcal{E}_{s}(\rho,\sigma,p)$ and gives their parameterization in terms of $(c,\psi_{\max},\Lambda_{\sigma})$, $(c,\psi_{\min},\Lambda_{\rho})$ or $(c,c_{r},\psi_{\max},\psi_{\min})$. The next Lemma evaluates acceptance for a given error-minimizing measurement as a function of these parameters $c$ and $c_{r}$.
Theorem 3.
For a test utilizing a measurement $\Lambda\in\mathcal{E}_{s}(\rho,\sigma,p)$, as given in Theorem 2, its acceptance under three cases is given as below.